# VIRAL: Vision-grounded Integration for Reward design And Learning

### v2

@@ Abstract
The refinement process can incorporate human feedback or be guided by a description generated by a video LLM, which explains the agent’s policy in video form. We evaluated VIRAL in five Gymnasium environments, demonstrating that it accelerates the learning of new behaviors while ensuring improved alignment with user intent. The source-code and demo video are available at: https://github.com/VIRAL-UCBL1/VIRAL and https://youtu.be/Hqo82CxVT38.
@@ 1 Introduction
We propose a Vision-grounded Integration for Reward design And Learning (VIRAL) framework to design rewards functions from simple users prompts and/or annotated image. VIRAL sets itself from the state-of-the-art [10, 12] in several key ways. First, VIRAL prioritizes the use of open-source, efficient, and lightweight LLMs [13, 14], ensuring greater accessibility, cost efficiency, and transparency. Second, unlike prior methods, VIRAL integrates Large Vision Language Models (LVLMs) [15, 16] to process both text and images, enhancing its ability to interpret user intent more accurately. Third, VIRAL is the first reward shaping approach to incorporate Video-LVLMs [17] for describing the movements of objects within the scene, providing richer context for reward function generation. Finally, instead of relying on direct access to the environment’s code (as in EUREKA [10]) or structured abstraction (through Pydantic class definitions as in Text2Reward [12]), VIRAL describes environments solely through its observations, following the Gymnasium [18] documentation. This simplifies implementation for users while ensuring the LLM captures the necessary information for coherent reward generation. Our main contributions are as follows: (1) The VIRAL pipeline to automatically design reward functions using a simple natural language prompt and/or annotated image. (2) A self-refinement process for reward functions, augmented with human feedback or video description from a LVLM. (3) A scalable and modular implementation designed to adapt to various RL problems within the Gymnasium framework, leveraging multiprocessing for faster inference. (4) An evaluation of VIRAL in five Gymnasium environments, showing its various benefits for learning and user intent alignment, using an empirical study.
@@ 2.1 The Input Parameters
To give the prompt $p_{g}$, the user has the possibility to choose between a text prompt, an annotated image $img$, or both of those. The $img$ can include annotations (with arrows, text, areas, etc.). The input $d_{env}$ is extracted from Gymnasium and provides an accurate representation of the environment’s observable space to the LLM, adding more depth to the generated reward function. Using text for $d_{env}$ allows for any Gymnasium environment to be seamlessly integrated with VIRAL. The input $f_{succ}:States(env)\to\{0,1\}$ is a success function which must be tailored to the goal. This function determines how success is manifested within the specific context of the task, helping the LLM in designing the reward function. Note that while $f_{succ}$ must be related to the goal prompt $p_{g}$, returning $1$ (success) can mean a partial success. For example, for a textual goal prompt $p_{g}:$ “Land the lander along the red trajectory shown in the image”, $f_{succ}$ may return $1$ if the lander managed to land, ignoring the trajectory.
@@ 2.3 Learning a Policy
The generated reward functions take an observation as parameters, along with two boolean values indicating whether there is a success or failure [^1] Note that an agent can be in a neutral state, without a success or a failure.. Once a reward function is generated, an RL algorithm is used to make the agent search for its policy. We restrict ourselves to Deep Q-Network (DQN) [21] and Proximal Policy Optimization (PPO) [22] and select the most suitable one for each environment (step 3 of Fig. 1). During training, rewards and observations are retrieved, and if the user wishes to, they can implement objective metric functions that take these observations and return a dictionary of useful objective metrics for this environment. During testing we can compare the success rate of our custom reward, with the baseline “legacy" reward function $r_{legacy}$ or with a user-defined threshold, which determines an acceptable success rate (step 4 of Fig. 1).
@@ 3 Empirical evaluation
For our evaluation, we used Qwen2.5-Coder-32B [23, 13] as the coder LLM due to its performance being comparable to that of GPT-4, making it a reliable choice for this role. For the multi-modal critic LLM, we chose Llama3.2-Vision-11B [16], as it is lightweight, open-source, and well-suited for our framework’s requirements. For Video-LVLM, we used Qwen2.5-VL-7B [24, 25] [^2] Our framework is model-agnostic and allows the use of any LLM for the coder and critic LLMs and Video-LVLM.. The evaluation was performed with a NVIDIA A40 GPU and a 12-cores Intel CPU.
@@
For the CartPole environment trained with PPO, and the goal prompt "Create a reward function for the CartPole environment that encourages keeping the pole upright for as long as possible.", Table 1 shows that the policy learned with our reward function outperforms that of Gymnasium. Additionally, we can see that the cart moves slightly more on average, resulting in a more stable pole overall.
Table 1: Comparison of legacy reward vs. ours (avg. over 10 runs)
@@
We conducted a study with 25 annotators to determine whether the learned behavior is aligned with the goal prompt provided (text, image or both) [^3] For a full description of the goal prompt used, we refer the reader to our Github repository.. Each annotator annotated a subset of 120 videos (4 environments and 10 videos per goal prompt modality). On average, each annotator rated 71 videos for a total of 1777 annotated videos. For each video, annotators used 5-point Likert scale (from “Strongly disagree” to “Strongly agree”) to answer the following two questions: (1) I understand the instructions, and (2) The instructions are followed. The human evaluation allowed us to assess the effectiveness of the modality on the semantic alignment.
For the first item, we obtained a mean of $4.89\pm 0.41$ and with only 147 videos with an understanding score less than 5/5, highlighting that very few goal prompts were misunderstood, confirming that goal prompts were well-crafted and understandable.
Fig. 2 shows the mean ratings for the second item and for each environment and modality. Overall, although using a visual goal prompt was not as good as a text-only goal prompt, we still obtained a mean of $2.14\pm 1.42$ (out of 5) across environments.
Figure 2: Semantic alignment for different environments and modalities.
However, as VIRAL’s main goal is to find the best reward, we investigated which modality led to very good behavior. Table 2 shows that, with the image modality, we discover behaviors that align more than with other modalities on Swimmer and Highway.
Table 2: Maximum selection of average rating per video, comparison by modality.
@@
Hopper-v5 3.41$\pm$1.00 4.92$\pm$0.28 4.83$\pm$0.39
LunarLander-v3 3.83$\pm$1.26 4.92$\pm$0.28 4.93$\pm$0.26
Swimmer-v5 4.19$\pm$0.75 4.14$\pm$0.95 3.41$\pm$1.33
Highway-fast-v0 4.44$\pm$1.29 4.06$\pm$1.03 3.91$\pm$1.38
These observations led us to the conclusion that different modalities can help us discover behaviors that align more closely with human intent, despite the number of correct behaviors in average.

### v3

@@ Abstract
The refinement process can incorporate human feedback or be guided by a description generated by a video LLM, which explains the agent’s policy in video form. We evaluated VIRAL in five Gymnasium environments, demonstrating that it accelerates the learning of new behaviors while ensuring improved alignment with user intent. The source-code and demo video are available at: https://github.com/VIRAL-UCBL1/VIRAL and https://youtu.be/BBIB4bEVSjE.
@@ 1 Introduction
We propose a Vision-grounded Integration for Reward design And Learning (VIRAL) framework to design rewards functions from simple users prompts and/or annotated image. VIRAL sets itself from the state-of-the-art [10, 12] in several key ways. First, VIRAL prioritizes the use of open-source, efficient, and lightweight LLMs [13, 14], ensuring greater accessibility, cost efficiency, and transparency. Second, unlike prior methods, VIRAL integrates Large Vision Language Models (LVLMs) [15, 16] to process both text and images, enhancing its ability to interpret user intent more accurately. Third, VIRAL is the first reward shaping approach to incorporate Video-LVLMs [17] for describing the movements of objects within the scene, providing richer context for reward function generation. Finally, instead of relying on direct access to the environment’s code (as in EUREKA [10]) or structured abstraction (through Pydantic class definitions as in Text2Reward [12]), VIRAL describes environments solely through its observations, following the Gymnasium [18] documentation. This simplifies implementation for users while ensuring the LLM captures the necessary information for coherent reward generation. Our main contributions are as follows:
  • The VIRAL pipeline to automatically design reward functions using a simple natural language prompt and/or annotated image.
  • A self-refinement process for reward functions, augmented with human feedback or video description from a LVLM.
  • A scalable and modular implementation designed to adapt to various RL problems within the Gymnasium framework, leveraging multiprocessing for faster inference.
  • An evaluation of VIRAL in five Gymnasium environments, showing its various benefits for learning and user intent alignment, using an empirical study.
@@ 2.1 The Input Parameters
To give the prompt $p_{g}$, the user has the possibility to choose between a text prompt, an annotated image imgimg, or both of those. The imgimg can include annotations (with arrows, text, areas, etc.). The input $d_{env}$ is extracted from Gymnasium and provides an accurate representation of the environment’s observable space to the LLM, adding more depth to the generated reward function. Using text for $d_{env}$ allows for any Gymnasium environment to be seamlessly integrated with VIRAL. The input $f_{succ}:States(env)\to\{0,1\}$ is a success function which must be tailored to the goal. This function determines how success is manifested within the specific context of the task, helping the LLM in designing the reward function. Note that while $f_{succ}$ must be related to the goal prompt $p_{g}$, returning $1$ (success) can mean a partial success. For example, for a textual goal prompt $p_{g}:$ “Land the lander along the red trajectory shown in the image”, $f_{succ}$ may return $1$ if the lander managed to land, ignoring the trajectory.
@@ 2.3 Learning a Policy
The generated reward functions take an observation as parameters, along with two boolean values indicating whether there is a success or failure (note that an agent can be in a neutral state, without a success or a failure). Once a reward function is generated, an RL algorithm is used to make the agent search for its policy. We restrict ourselves to Deep Q-Network (DQN) [21] and Proximal Policy Optimization (PPO) [22] and select the most suitable one for each environment (step 3 of Fig. 1). During training, rewards and observations are retrieved, and if the user wishes to, they can implement objective metric functions that take these observations and return a dictionary of useful objective metrics for this environment. During testing we can compare the success rate of our custom reward, with the baseline “legacy" reward function $r_{legacy}$ or with a user-defined threshold, which determines an acceptable success rate (step 4 of Fig. 1).
@@ 3 Empirical evaluation
For our evaluation, we used Qwen2.5-Coder-32B [23, 13] as the coder LLM due to its performance being comparable to that of GPT-4, making it a reliable choice for this role. For the multi-modal critic LLM, we chose Llama3.2-Vision-11B [16], as it is lightweight, open-source, and well-suited for our framework’s requirements. For Video-LVLM, we used Qwen2.5-VL-7B [24, 25]. However, we note that our framework is model-agnostic and allows the use of any LLM for the coder and critic LLMs and Video-LVLM. The evaluation was performed with a NVIDIA A40 GPU and a 12-cores Intel CPU.
@@
For the CartPole environment trained with PPO, and the goal prompt "Create a reward function for the CartPole environment that encourages keeping the pole upright for as long as possible.", Table 1 shows that the policy learned with our reward function outperforms that of Gymnasium. We limited the number of epochs to 25 000 in order to make the learning problem more complex and suitable for evaluating learning speed. Additionally, we can see that the cart moves slightly more on average, resulting in a more stable pole overall.
Table 1: Comparison of legacy reward vs. ours on Cartpole Environment(avg. over 10 runs)
@@
We conducted a study with 25 annotators to determine whether the learned behavior is aligned with the goal prompt provided (text, image or both). For a full description of the goal prompt used, we refer the reader to our Github repository. Each annotator was tasked to annotate a maximum of 120 videos (4 environments and 10 videos per goal prompt modality). On average, each annotator rated 71 videos for a total of 1777 annotated videos. For each video, annotators used 5-point Likert scale (from “Strongly disagree” to “Strongly agree”) to answer the following two questions: (1) I understand the instructions, and (2) The instructions are followed. The human evaluation allowed us to assess the effectiveness of the modality on the semantic alignment.
For the first item, we obtained a mean of $4.89\pm 0.41$ and with only 147 (out of 1777) videos with an understanding score less than 5/5, highlighting that very few goal prompts were misunderstood, confirming that goal prompts were well-crafted and understandable. This indicates that the study’s results are therefore not confounded by ambiguity in the prompts themselves.
Fig. 2 shows the mean ratings for the second item (over the 10 videos of each annotator) for each environment and modality. We note that text-only prompts resulted in the highest semantic alignment, indicating that textual descriptions provided clearer, more actionable guidance for the system than other modalities. Overall, although using an image-only prompt was not as good as a text-only goal prompt, we still obtained a mean of $2.14\pm 1.42$ (out of 5) across environments for the image modality. Adding text to image goal prompts did not consistently improve alignment, and in some cases hurt performance (increase in Hopper and LunarLander environments and decrease in the Swimmer and Highway-fast environments), suggesting potential interference or confusion introduced by multimodal inputs. The relative performance of each modality varied with the environment, which implies that the usefulness of multimodal or visual instructions may depend heavily on the task’s nature.
Figure 2: Semantic alignment over the 10 videos of each annotator for different gymnasium’s environments and modalities.
However, we emphasize that VIRAL’s primary objective is to discover the optimal reward across multiple generations. Therefore, we argue that maximizing semantic alignment in the best-performing instances is more relevant than focusing solely on the average alignment across all videos for a given modality. Specifically, for each video, we compute the average semantic alignment score reported by the annotators, and then, for each modality, we report the highest of these averages across the 10 videos in Table 2. We can make the following three observations: (1) Textual goal prompts generally lead to the best alignment. Indeed, in the Hopper environment, the textual modality achieves the highest score $(4.92\pm 0.28)$, outperforming both image and image+text. Moreover, in the other environments, the textual modality is nearly tied to the best modality. This suggests that textual goal prompts alone are highly effective at conveying the intended behavior across diverse environments. (2) Image-only goal prompts can occasionally perform best (in the Swimmer and Highway-fast environments), but are less consistent. This indicates that visual cues alone can be sufficient, but their effectiveness is highly environment-dependent — likely tied to how interpretable or informative the visual goal is for that task. Moreover, the variance is relatively high which could imply greater inconsistency in human judgments when interpreting the alignment with image-only goal prompts. (3) Text+image goal prompts can improve alignment but does not improve it consistently. While image+text performs well in LunarLander ($4.93\pm 0.26$), it underperforms in the other modalities. This reinforces our previous finding that adding text to an image goal prompt does not always help, and might even introduce noise or ambiguity in some environments.
Table 2: Maximum Average Semantic Alignment per Modality and Environment.
@@
[…]